OpenStudy (anonymous):

A ball is thrown horizontally from the roof of a building 57 m tall and lands 40 m from the base. What was the ball's initial speed?

OpenStudy (souvik):

let v be the initial horizontal velocity.. suppose after time t the ball lands then \[vt=40\] ............(1) the initial vertical velocity of the ball is zero we can use this equation...\[h _{y}=u _{y}t+1/2g t ^{2}\] here \[u _{y}=0\] and \[h _{y}=57\] then \[57=1/2g t ^{2}\]..................(2) eliminate t from this two equation..you will get the anser.. \[v=11.7 m/s\]

OpenStudy (anonymous):

THanks

OpenStudy (souvik):

welcome

OpenStudy (souvik):

if done close the question

OpenStudy (anonymous):

An athlete executing a long jump leaves the ground at a 35° angle and travels 6.80 m. ) What was the takeoff speed? (b) If this speed were increased by just 5.0%, how much longer would the jump be?

OpenStudy (souvik):

hint:draw a simple picture..let v be the velocity and t be the time..divide the velocity into two components...make two equation and find the unknown v by eliminating t..try it

OpenStudy (anonymous):

when you say two components do you mean in the x direction and they direction

OpenStudy (souvik):

take a component horizontally (X component) another vertically(Y component)

OpenStudy (anonymous):

but in order to find those, dont i need the velocity

OpenStudy (souvik):

let v be the velocity..the X component of the velocity is\[v \cos \theta\] and Y component is \[v \sin \theta \] set two equation...and solve it

OpenStudy (anonymous):

You know that in the x direction you traveled 6.80 m right? So \[v \cos \theta t = 6.80 m\] We also know that the only acceleration in the y direction is gravity. The motion look like this.  The final velocity in the y direction is 0 right? So you need to substitute t with \[vsin \theta \] and solve for takeoff speed.